the perimeter of a rectangle is 180m the length is 5m more than double the width use an equation
Answer by mananth(16946)

the perimeter of a rectangle is 180m the length is 5m more than double the width use an equation

width=x
length=2x+5
2*(l+w)=perimeter
2*(x+2x+5)=180
/2
3x+5=90
-5 -5
3x=85
/3
x=28.33m the width
length = 2x+5 =61.66
...
m.ananth@hotmail.ca